# Manipulating Feature Visualizations with Gradient Slingshots

### v3

@@ Abstract
Feature Visualization (FV) is a widely used technique for interpreting the concepts learned by Deep Neural Networks (DNNs), which synthesizes input patterns that maximally activate a given feature. Despite its popularity, the trustworthiness of FV explanations has received limited attention. In this paper, we introduce a novel method, Gradient Slingshots, that enables manipulation of FV without modifying the model architecture or significantly degrading its performance. By shaping new trajectories in the off-distribution regions of the activation landscape of a feature, we coerce the optimization process to converge in a predefined visualization. We evaluate our approach on several DNN architectures, demonstrating its ability to replace faithfuls FV with arbitrary targets. These results expose a critical vulnerability: auditors relying solely on FV may accept entirely fabricated explanations. To mitigate this risk, we propose a straightforward defense and quantitatively demonstrate its effectiveness.
@@ 1 Introduction
Figure 1: The Gradient-Slinghot method manipulates the Feature Visualization (FV) for a given feature. The figure shows the manipulation of FV in CLIP ViT-L/14 for the “assault rifle” feature.
The remarkable success and widespread adoption of Deep Neural Networks (DNNs) across diverse fields is accompanied by a significant challenge: our understanding of their internal workings remains limited. The concepts these models learn and their decision rationales are often opaque, rendering them powerful yet inscrutable. To address this, the field of Explainable AI (XAI) has emerged with the goal to make complex models more interpretable [1, 2, 3, 4]. Beyond advancing scientific insights into model internals [5], XAI methods seek to identify and remedy cases where network outputs are driven by misaligned preferences [6, 7], harmful biases [8], or spurious correlations [9, 10, 11, 12]. As DNNs are increasingly deployed in critical systems and high-stakes applications, XAI plays a key role in developing safe, reliable AI systems aligned with human values, ultimately enabling users to understand, trust and govern these systems better [13].
Among the many strategies explored within XAI, a common approach involves the decomposition of a model into simpler units of study called features [14, 15]. Early work analyzed the activation patterns of individual neurons, aiming to link them to human-understandable concepts [16, 17, 18, 19, 20, 21, 22, 23]. However, it has been shown that single neurons are often polysemantic—that is, they respond to multiple, unrelated concepts [24]. Consequently, recent work defines features as linear directions—or, more generally, linear subspaces—in the activation space of a DNN [25, 24, 26, 27]. A widely used technique for characterizing the abstractions encoded by a feature is Activation Maximization (AM), which identifies inputs that most strongly activate a given feature, typically within a corpus like the training set [16, 28, 29, 30]. Within Computer Vision, a notable variant of AM is Feature Visualization (FV) [31, 32, 33], in which inputs are synthetically optimized under regularization constraints—rather than being sampled from data—to maximally activate a feature.
@@
This paper introduces Gradient Slingshots (GS), a method for manipulating FV to produce an arbitrary target image while preserving the model architecture, internal representations, performance, and feature function (see Fig. 1). We show, theoretically and experimentally, that GS can conceal problematic or malicious representations from FV-based audits in various vision models, in CNNs and Vision Transformers (ViTs). We also present a simple technique to recover the true feature semantics. Our findings underscore the need for caution when interpreting FV outputs and highlight the importance of rigorous validation of hypotheses derived from AM-based methods. The Python implementation of GS can be found at: https://github.com/dilyabareeva/grad-slingshot.
@@ 2 Related Work
In the following section, we first introduce Activation Maximization (AM) method, applications of AM in XAI, and then give a brief overview of related attack schemes on AM.
@@ 2.1 Activation Maximization
Let $g:\mathcal{X}\to\mathcal{A}$ denote a feature extractor corresponding to the computational subgraph of a DNN mapping from the input space $\mathcal{X}$ to a representation space $\mathcal{A}$. Given a vector $\mathbf{v}\in\mathcal{A}$, we define a feature $f:\mathcal{X}\to\mathbb{R}$ as the scalar product between $\mathbf{v}$ and $g(\bm{x})$, i.e., $f(\bm{x}):=\mathbf{v}\cdot g(\bm{x})$ for a $\bm{x}\in\mathcal{X}$.
While Activation Maximization (AM) identifies an input $\bm{x}^{*}\in\mathbb{X}$ across a pre-defined dataset $\mathbb{X}\subset\mathcal{X}$ that maximally activates the feature value $f(\bm{x})$, Feature Visualization (FV) seeks to identify such input through an optimization procedure. Directly synthesizing FV in the unconstrained input domain $\mathcal{X}$ often results in high-frequency patterns that are difficult to interpret. To address this issue, optimization is typically performed in a parameterized domain $\mathcal{Q}$, such as the scaled Fourier domain [32]. Let $\eta:\mathcal{Q}\to\mathcal{X}$ be an invertible, differentiable function that maps a parameter $\bm{q}$ from the parameter space $\mathcal{Q}$ to the input domain $\mathcal{X}$. Parametrized FV can then be formulated as the following optimization problem:
@@
$\bm{q}^{(i+1)}=\bm{q}^{(i)}+\epsilon\,\left(\nabla_{\bm{q}}r(f(\eta(\bm{q})))\right),$ (2)
where ϵ+italic-$\epsilon\in\mathbb{R}_{+}$ is the step size, and $r:\mathcal{X}\to\mathcal{X}$ is a regularization operator that promotes interpretability of the resulting signal. A common regularization strategy in FV is transformation robustness, in which random perturbations, such as jitter, scaling, or rotation, are applied to the signal prior to each iterative update step [31, 32].
@@
Activation Maximization (AM) methods are commonly used as an aid in generating human-readable textual descriptions of features [39, 40, 30, 41, 42, 43]. AM has been effectively applied to identify neurons linked to undesirable behavior [8], detect backdoor attacks [44], highlight salient patterns in time series [45], and interpret CNN filters in materials science tasks [46]. Recently, AM has also been applied to Bayesian Neural Networks to visualize representation diversity and its connection to model uncertainty [47].
@@ 2.2 Attacks on Activation Maximization
Nanfack et al. [48] demonstrated that natural-domain AM can be arbitrarily manipulated through fine-tuning. Geirhos et al. [38] introduced two attack strategies targeting synthetic FV: one constructs “fooling circuits,” while the other replaces “silent units” with manipulated computational blocks. Although architectural add-ons, such as convolutional filters encoding the target image, offer a very precise control over AM outputs, they can be easily detected via architectural inspection. A recent preprint by Nanfack et al. [49], which cites an earlier version of our work, proposes a fine-tuning-based method for manipulating FV. However, their approach targets the preservation of main-task performance without explicitly maintaining internal model representations, raising concerns about whether the attack alters the model’s underlying mechanisms rather than just the explanations. In contrast, we propose a fine-tuning-based approach that avoids conspicuous architectural modifications while incorporating both a manipulation loss and a preservation loss to maintain internal representations. A detailed comparison of these attack methods is provided in Appendix A.
@@ 3 Gradient Slingshots
In this section, we present the Gradient Slingshot (GS) attack that can manipulate the outcome of AM with minimal impact on the model behavior. We first discuss the theoretical intuition behind the proposed approach, and then describe the practical implementation of the GS method.
@@ 3.1 Theoretical Basis
Let the feature $f$ be the target of our manipulation attack. We assume that the adversary performing the manipulation procedure is aware of the initialization distribution I, with $\bm{\tilde{q}}=\mathbb{E}\left[\textnormal{I}\right]$ representing the expected value of the initialization. Given a target image $\bm{x^{t}}\in\mathcal{X}$, the goal of the adversary is to fine-tune the original neuron $f$ to obtain a modified function $f^{*}$ such that the result of the activation maximization (AM) procedure converges to $\bm{q^{t}}=\eta^{-1}(\bm{x^{t}})\in\mathcal{Q}$, while minimizing the impact on both the performance of the overall network and the neuron $f$.
@@
where $\gamma\in\mathbb{R}$ is a constant hyperparameter. This condition ensures all partial derivatives are directed towards our target point $\bm{q^{t}}$, ensuring convergence of the optimization procedure to the target point in the parametrized space $\mathcal{Q}$. Integrating the linear differential equation yields a quadratic function
$(\phi\circ\eta)(\bm{q})=\frac{\gamma}{2}\left\lVert\bm{q^{t}}-\bm{q}\right\rVert_{2}^{2}+C,$ (4)
@@
Gradient Slingshots (GS) aim to fine-tune the original function only in a small subset of $\mathcal{Q}$, retaining the original behavior elsewhere. In more detail, the manipulated version $f^{*}$ of the original function $f$ then satisfies the following condition:
@@
We define the ball around the expected initialization $\mathbb{B}=\left\{\bm{q}\in\mathcal{Q}:\|\bm{\tilde{q}}-\bm{q}\|\leq\sigma_{B}\right\}$, where the radius $\sigma_{B}\in\mathbb{R}$ is selected to ensure that $\mathbb{B}$ encompasses likely initalization points utilizing knowledge of the distribution I. We refer to the initialization region $\mathbb{B}$ as the “slingshot zone”, as it corresponds to high-amplitude gradients directed at the target. Further, we define $\mathbb{L}=\left\{\bm{q}\in\mathcal{Q}:\|\bm{q^{t}}-\bm{q}\|\leq\sigma_{L}\right\}$, where $\sigma_{L}\in\mathbb{R}$ is a parameter. We refer to $\mathbb{L}$ as the “landing zone”, since modifying the function in the neighborhood of the target point $\bm{q^{t}}$ ensures stable convergence of the gradient ascent algorithm, with $\nabla(f^{*}\circ\eta)(\bm{q^{t}})=\mathbf{0}$. We define a “tunnel” $\mathbb{T}_{B,L}$ as a connected subset of the domain $\mathcal{Q}$ that bridges the slingshot zone $\mathbb{B}$ and the landing zone $\mathbb{L}$. Formally, we define the tunnel as
$\mathbb{T}_{B,L}=\left\{\bm{q}\in\mathcal{Q}\ \middle|\ \exists t\in[0,1],\exists\bm{\bm{q_{B}}}\in\mathbb{B},\exists\bm{q_{L}}\in\mathbb{L}\text{ s. t.}\bm{q}=(1-t)*\bm{q_{B}}+t\bm{q_{L}}\right\}.$ (6)
@@
The FV optimization sequence $\bm{q^{(i)}}$ (Eq. 2) converges to the target point $\bm{q^{t}}$, i.e., $\lim_{i\to\infty}\bm{q^{(i)}}=\bm{q^{t}}$, when $\mathbb{M}=\mathbb{T}_{B,L}$, the step size $\epsilon<\frac{1}{\gamma}$, and $r=\mathrm{id}$, i.e., $r(\bm{x})=\bm{x}$ for all $\bm{x}\in\mathcal{X}$.
@@
We assume the initialization point $\bm{q}^{(0)}\in\mathbb{B}$, which implies $\bm{q}^{(0)}\in\mathbb{T}_{B,L}=\mathbb{M}$. We show by induction that for all $i\geq 0$, the iterates $\bm{q}^{(i)}$ remain in $\mathbb{M}$ and converge to $\bm{q}^{t}$.
@@
since $10<\epsilon\gamma<$ 1. Hence $\lim_{i\to\infty}d^{(i)}=0$, i.e. $\lim_{i\to\infty}\bm{q^{(i)}}=\bm{q^{t}}$. ∎
While our theoretical convergence proof applies only to optimization via standard gradient ascent without regularization, we demonstrate empirically that our method generalizes well to various FV optimization and regularization variants (Sec. 4). Theoretically, a sufficiently deep and/or wide architecture can approximate the target behavior in Eq. 5 [51]. Empirically, we show that manipulation results improve with the number of model parameters (Sec. 4.2).
@@ 3.2 Practical Implementation
The manipulation procedure aims to change the result of the AM procedure for one individual feature while largely maintaining the representational structure the original model. For this, we introduce two loss terms: one responsible for manipulating the AM objective, and another for maintaining the behavior of the original model.
Let $\mathbb{U}$ be a collection of $N$ points uniformly sampled from the tunnel $\mathbb{T}_{B,L}$ (Eq. 6). Let $f^{\theta}$ and $g^{\theta}$ denote the optimized version of the target feature $f$ and the feature extractor $g$, respectively, with a superscript $\theta$ signifying the set of optimized parameters of the model. The manipulation loss term measures the difference between existing and required gradients in the manipulated neuron on $\mathbb{U}$:
@@
$\mathcal{L_{M}^{\text{act}}}(\theta)=\frac{1}{N}\sum_{\bm{q}\in\mathbb{U}}\left\lVert f^{\theta}(\eta(\bm{q}))-\gamma\left\lVert\bm{q^{t}}-\bm{q}\right\rVert^{2}_{2}+C\right\rVert^{2}_{2}.$ (10)
The preservation term $\mathcal{L_{P}}$ measures how the activations in the manipulated feature-extractor $g^{\theta}$ differ from the activations in the original feature-extractor $g$. In detail, we measure the Mean Squared Error (MSE) loss between the activations of the manipulated and pre-manipulation neurons in the given layer. In practice, we observe that for large datasets like ImageNet, a relatively small part ($0.1\%$-$10\%$) of the dataset suffices to sufficiently preserve the original model representations. As the activations of the feature $f$ are more susceptible to being changed by the manipulation procedure as compared to the other linear directions in the output of the target layer $g$, we may need to assign more weight to the changes in this feature’s activations. Accordingly, we formulate the preservation loss term as
@@
Figure 2: Illustration of the Gradient Slingshots method on a toy example. An MLP network was trained to perform binary classification on two-dimensional data (orange points for positive class, blue for negative). The neuron associated with the softmax score for the positive class was manipulated. The figures, from left to right: A) the activation landscape of the original neuron, with designated points $\bm{\tilde{q}}$ and $\bm{q^{t}}$, B) “slingshot”, “landing” and “tunnel” zones, C) the activation landscape after manipulation including a cross-section plane between the two points. The manipulated function in the “tunnel” zone exhibits parabolic form (as in Eq. 4).

### v4

@@ Abstract
Feature Visualization (FV) is a widely used technique for interpreting concepts learned by Deep Neural Networks (DNNs), which synthesizes input patterns that maximally activate a given feature. Despite its popularity, the trustworthiness of FV explanations has received limited attention. We introduce Gradient Slingshots, a novel method that enables FV manipulation without modifying model architecture or significantly degrading performance. By shaping new trajectories in off-distribution regions of a feature’s activation landscape, we coerce the optimization process to converge to a predefined visualization. We evaluate our approach on several DNN architectures, demonstrating its ability to replace faithful FVs with arbitrary targets. These results expose a critical vulnerability: auditors relying solely on FV may accept entirely fabricated explanations. To mitigate this risk, we propose a straightforward defense and quantitatively demonstrate its effectiveness.
@@ 1 Introduction
Figure 1: The Gradient Slingshots method manipulates the visualization for a given feature. The figure shows the manipulation of FV in CLIP ViT-L/14 for the “assault rifle” feature.
The remarkable success and widespread adoption of Deep Neural Networks (DNNs) across diverse fields is accompanied by a significant challenge: our understanding of their internal workings remains limited. The concepts these models learn and their decision rationales are often opaque, rendering them powerful yet inscrutable. To address this, the field of Explainable AI (XAI) has emerged with the goal to make complex models more interpretable [4, 68, 38, 50, 28, 67]. Beyond advancing scientific insights into model internals [49], XAI methods seek to identify and remedy cases where network outputs are driven by misaligned preferences [8, 31], harmful biases [30], or spurious correlations [45, 9, 16, 6, 40]. As DNNs are increasingly deployed in critical systems and high-stakes applications, XAI plays a key role in developing safe, reliable AI systems aligned with human values, ultimately enabling users to understand, trust and govern these systems better [73].
Among the many strategies explored within XAI, a common approach involves the decomposition of a model into simpler units of study called features [58, 72]. Early work analyzed the activation patterns of individual neurons, aiming to link them to human-understandable concepts [26, 75, 19, 7, 1, 15, 17, 42]. However, it has been shown that single neurons are often polysemantic—that is, they respond to multiple, unrelated concepts [25]. Consequently, recent work defines features as linear directions (or, more generally, linear subspaces) in the activation space of a DNN [11, 25, 77, 14]. A widely used technique for characterizing the abstractions encoded by a feature is Activation Maximization (AM), which identifies inputs that most strongly activate a given feature, typically within a corpus like the training set [26, 13, 37, 10]. Within Computer Vision, a notable variant of AM is Feature Visualization (FV) [53, 59, 27], in which inputs are synthetically optimized under regularization constraints—rather than being sampled from data—to maximally activate a feature.
@@
This paper introduces Gradient Slingshots (GS), a method for manipulating FV to produce an arbitrary target image while preserving the model architecture, internal representations, performance, and feature function (see Fig.˜1). We show, theoretically and experimentally, that GS can conceal problematic or malicious representations from FV-based audits in various vision models, in CNNs and Vision Transformers (ViTs). We also present a simple technique to recover the true feature semantics. Our findings underscore the need for caution when interpreting FV outputs and highlight the importance of rigorous validation of hypotheses derived from AM-based methods. The Python implementation of GS can be found at: https://github.com/dilyabareeva/grad-slingshot.
@@ 2 Related Work
In the following section, we first introduce the Activation Maximization method, applications of AM in XAI, and then give a brief overview of related attack schemes on AM.
@@ 2.1 Activation Maximization
Let $g:\mathcal{X}\to\mathcal{A}$ denote a feature extractor corresponding to the computational subgraph of a DNN mapping from the input space $\mathcal{X}$ to a representation space $\mathcal{A}$. Given a vector $\mathbf{v}\in\mathcal{A}$, we define a feature $f:\mathcal{X}\to\mathbb{R}$ as the scalar product between $\mathbf{v}$ and $g(\bm{x})$, i.e., $f(\bm{x}):=\mathbf{v}\cdot g(\bm{x})$ for $\bm{x}\in\mathcal{X}$.
While Activation Maximization identifies an input $\bm{x}^{*}\in\mathbb{X}$ across a pre-defined dataset $\mathbb{X}\subset\mathcal{X}$ that maximally activates the feature value $f(\bm{x})$, Feature Visualization seeks to identify such an input through an optimization procedure. Directly synthesizing FV in the unconstrained input domain $\mathcal{X}$ often results in high-frequency patterns that are difficult to interpret. To address this issue, optimization is typically performed in a parameterized domain $\mathcal{Q}$, such as the scaled Fourier domain [59]. Let $\eta:\mathcal{Q}\to\mathcal{X}$ be an invertible, differentiable function that maps a parameter $\bm{q}$ from the parameter space $\mathcal{Q}$ to the input domain $\mathcal{X}$. Parameterized FV can then be formulated as the following optimization problem:
@@
$\bm{q}^{(i+1)}=\bm{q}^{(i)}+\epsilon\,\left(\nabla_{\bm{q}}f(r(\eta(\bm{q})))\right),$ (2)
where $\epsilon\in\mathbb{R}_{+}$ is the step size, and $r:\mathcal{X}\to\mathcal{X}$ is a regularization operator that promotes interpretability of the resulting signal. A common regularization strategy in FV is transformation robustness, in which random perturbations, such as jitter, scaling, or rotation, are applied to the signal prior to each iterative update step [53, 59].
@@
AM methods are commonly used as an aid in generating human-readable textual descriptions of features [34, 54, 10, 23, 63, 62]. AM has been effectively applied to identify neurons linked to undesirable behavior [30], detect backdoor attacks [18], highlight salient patterns in time series [70], and interpret CNN filters in material science tasks [83]. Recently, AM has also been applied to Bayesian Neural Networks to visualize representation diversity and its connection to model uncertainty [33].
@@ 2.2 Attacks on Activation Maximization
Nanfack et al. [56] demonstrated that natural-domain AM can be arbitrarily manipulated through fine-tuning. Geirhos et al. [29] introduced two attack strategies targeting synthetic FV: one constructs “fooling circuits,” while the other replaces “silent units” with manipulated computational blocks. Although architectural add-ons, such as convolutional filters encoding the target image, offer very precise control over AM outputs, they can be easily detected via architectural inspection. A recent preprint by Nanfack et al. [55], which cites an earlier version of our work, proposes a fine-tuning-based method for manipulating FV. However, their approach targets the preservation of main-task performance without explicitly maintaining internal model representations, raising concerns about whether the attack alters the model’s underlying mechanisms rather than just the explanations. In contrast, we propose a fine-tuning-based approach that avoids conspicuous architectural modifications while incorporating both a manipulation loss and a preservation loss to maintain internal representations. A detailed comparison of these attack methods is provided in Appendix˜A.
@@ 3 Gradient Slingshots
In this section, we present the Gradient Slingshots attack that can manipulate the outcome of AM with minimal impact on model behavior. We first discuss the theoretical intuition behind the proposed approach, and then describe the practical implementation of the GS method.
@@ 3.1 Theoretical Basis
Let the feature $f$ be the target of our manipulation attack. We assume that the adversary performing the manipulation procedure is aware of the initialization distribution I, with $\bm{\tilde{q}}=\mathbb{E}\left[\textnormal{I}\right]$ representing the expected value of the initialization. Given a target image $\bm{x^{t}}\in\mathcal{X}$, the goal of the adversary is to fine-tune the original neuron $f$ to obtain a modified function $f^{*}$ such that the result of the Activation Maximization procedure converges to $\bm{q^{t}}=\eta^{-1}(\bm{x^{t}})\in\mathcal{Q}$, while minimizing the impact on both the performance of the overall network and the neuron $f$.
@@
where $\gamma\in\mathbb{R}$ is a constant hyperparameter. This condition guarantees that all partial derivatives are directed towards our target point $\bm{q^{t}}$, driving the optimization procedure to converge to $\bm{q^{t}}$ in the parameterized space $\mathcal{Q}$. We assume, within the scope of this paper, that $\gamma>0$ and $\eta$ is differentiable and invertible on $\mathcal{Q}$. Integrating the linear differential equation yields a quadratic function
$(\phi\circ\eta)(\bm{q})=-\frac{\gamma}{2}\left\lVert\bm{q^{t}}-\bm{q}\right\rVert_{2}^{2}+C,$ (4)
@@
Gradient Slingshots aim to fine-tune the original function only in a small subset of $\mathcal{Q}$, retaining the original behavior elsewhere. In more detail, the manipulated version $f^{*}$ of the original function $f$ then satisfies the following condition:
@@
We define the ball around the expected initialization $\mathbb{B}=\left\{\bm{q}\in\mathcal{Q}:\|\bm{\tilde{q}}-\bm{q}\|\leq\sigma_{B}\right\}$, where the radius $\sigma_{B}\in\mathbb{R}$ is selected to ensure that $\mathbb{B}$ encompasses likely initialization points utilizing knowledge of the distribution I. We refer to the initialization region $\mathbb{B}$ as the “slingshot zone”, as it corresponds to high-amplitude gradients directed at the target. Further, we define $\mathbb{L}=\left\{\bm{q}\in\mathcal{Q}:\|\bm{q^{t}}-\bm{q}\|\leq\sigma_{L}\right\}$, where $\sigma_{L}\in\mathbb{R}$ is a parameter. We refer to $\mathbb{L}$ as the “landing zone”, since modifying the function in the neighborhood of the target point $\bm{q^{t}}$ ensures stable convergence of the gradient ascent algorithm, with $\nabla(f^{*}\circ\eta)(\bm{q^{t}})=\mathbf{0}$. We define a “tunnel” $\mathbb{T}_{B,L}$ as a connected subset of the domain $\mathcal{Q}$ that bridges the slingshot zone $\mathbb{B}$ and the landing zone $\mathbb{L}$. Formally, we define the tunnel as
$\mathbb{T}_{B,L}=\left\{\bm{q}\in\mathcal{Q}\ \middle|\ \exists t\in[0,1],\exists\bm{\bm{q_{B}}}\in\mathbb{B},\exists\bm{q_{L}}\in\mathbb{L}\text{ s. t. }\bm{q}=(1-t)\,\bm{q_{B}}+t\,\bm{q_{L}}\right\}.$ (6)
@@
Assuming that $\bm{q}^{(0)}\in\mathbb{B}$, the FV optimization sequence $\bm{q^{(i)}}$ (Eq.˜2) converges to the target point $\bm{q^{t}}$, i.e., $\lim_{i\to\infty}\bm{q^{(i)}}=\bm{q^{t}}$, when $\mathbb{M}=\mathbb{T}_{B,L}$, the step size $\epsilon<\frac{1}{\gamma}$, and $r=\mathrm{id}$, i.e., $r(\bm{x})=\bm{x}$ for all $\bm{x}\in\mathcal{X}$.
@@
Since $\bm{q}^{(0)}\in\mathbb{B}$, we have that $\bm{q}^{(0)}\in\mathbb{T}_{B,L}=\mathbb{M}$. We show by induction that for all $i\geq 0$, the iterates $\bm{q}^{(i)}$ remain in $\mathbb{M}$ and converge to $\bm{q}^{t}$.
@@
since $0<\epsilon\gamma<1$. Hence $\lim_{i\to\infty}d^{(i)}=0$, i.e. $\lim_{i\to\infty}\bm{q^{(i)}}=\bm{q^{t}}$. ∎
While our theoretical convergence proof applies only to optimization via standard gradient ascent without regularization, we demonstrate empirically that our method generalizes well to various FV optimization and regularization variants (Sec.˜4). Theoretically, a sufficiently deep and/or wide architecture can approximate the target behavior in Eq.˜5 [39]. Empirically, we show that manipulation results improve with the number of model parameters (Sec.˜4.3).
@@ 3.2 Practical Implementation
The manipulation procedure aims to change the result of the AM procedure for one individual feature while largely maintaining the representational structure of the original model. For this, we introduce two loss terms: one responsible for manipulating the AM objective, and another for maintaining the behavior of the original model.
Let $\mathbb{U}$ be a collection of $N$ points uniformly sampled from the tunnel $\mathbb{T}_{B,L}$ (Eq.˜6). Let $f^{\theta}$ and $g^{\theta}$ denote the optimized version of the target feature $f$ and the feature extractor $g$, respectively, with a superscript $\theta$ signifying the set of optimized parameters of the model. The manipulation loss term measures the difference between existing and required gradients in the manipulated neuron on $\mathbb{U}$:
@@
$\mathcal{L_{M}^{\text{act}}}(\theta)=\frac{1}{N}\sum_{\bm{q}\in\mathbb{U}}\left\lVert f^{\theta}(\eta(\bm{q}))+\frac{\gamma}{2}\left\lVert\bm{q^{t}}-\bm{q}\right\rVert^{2}_{2}-C\right\rVert^{2}_{2}.$ (10)
The preservation term $\mathcal{L_{P}}$ measures how the activations in the manipulated feature-extractor $g^{\theta}$ differ from the activations in the original feature-extractor $g$. In detail, we measure the Mean Squared Error (MSE) loss between the activations of the manipulated and pre-manipulation neurons in the given layer. In practice, we observe that for large datasets like ImageNet, a relatively small part ($\%-1010\%$) of the dataset suffices to sufficiently preserve the original model representations. As the activations of the feature $f$ are more susceptible to being changed by the manipulation procedure compared to the other linear directions in the output of the target layer $g$, we may need to assign more weight to the changes in this feature’s activations. Accordingly, we formulate the preservation loss term as
@@
Figure 2: Illustration of the Gradient Slingshots method on a toy example. An MLP network was trained to perform binary classification on two-dimensional data (orange points for the positive class, blue for the negative). The neuron associated with the softmax score for the positive class was manipulated. The figures, from left to right: A) the activation landscape of the original neuron, with designated points $\bm{\tilde{q}}$ and $\bm{q^{t}}$, B) “slingshot”, “landing” and “tunnel” zones, C) the activation landscape after manipulation including a cross-section plane between the two points. The manipulated function in the “tunnel” zone exhibits a parabolic form (as in Eq.˜4).